OpenStudy (anonymous):

solve the multiple-angle equation cos x/4=0

OpenStudy (anonymous):

\[\cos(4x)=0\\ 4x=\frac{\pi}{2}\]

OpenStudy (anonymous):

how did he get this? x/4 = ± π/2 + 2nπ